Calculate the forces in members BE and BD of the loaded truss. The forces are positive if in tension, negative if in compression. Assume F = 4.4 kN, a = 2.3 m, b = 3 m, c = 2 m, and d = 2 m.


1
Expert's answer
2021-08-04T02:51:54-0400

In order to find BE we will resolve force at point E

"\sum F_y=0\\
 F_{BE}=0"

Force in member BE = 0

In order to find force in BD, we can resolve force along a vertical direction at point B

"\sum F_y=0\\
 F_{BE}+F_{AB} cos \theta+F_{BD} cos \theta=0\\
Since \space F_{BE} =0\\
F_{AB} cos \theta = -F_{BD} cos \theta \\
F_{AB}=-F_{BD}-------(1)"

Now we need to find FAB

At point A resolving along the vertical direction

"F= F_{AB}sin \theta\\
F= F_{AB} \frac{3}{\sqrt{13}}\\ 
F_{AB}= \frac{\sqrt{13}}{3}(4.4)\\ 
F= 5.28814kN (Tension)"

Substituting in equation 1

"F_{AB}= -F_{BD}\\
5.28814= -F_{BD}\\
F_{BD}=-5.28814(Compression)"